# An Algorithmic Approach to Finding Degree-Doubling Nodes in Oriented Graphs

### v1

@@ Abstract
Seymour’s Second Neighborhood Conjecture asserts that in the square of any oriented graph, there exists a node whose out-degree at least doubles. This paper presents a definitive proof of the conjecture by introducing the GLOVER (Graph-Level Oriented Vertex Expansion and Reduction) data structure, which facilitates a systematic partitioning of neighborhoods and an analysis of degree-doubling conditions. By leveraging this structure, we construct a decreasing sequence of subsets that establish a well-ordering of nodes, ensuring that no counterexample can exist. This approach not only confirms the conjecture for all oriented graphs but also provides a novel framework for analyzing degrees and arcs in complex networks. The findings have implications for theoretical graph studies and practical applications in network optimization and algorithm design.
@@ A Graph Level Order (GLOVER) on set $S$ can be defined as follows:
Cycle in Container 1 If $G$ is an oriented graph represented as a GLOVER with $v_{0}$ as the chosen minimum degree node, and if there exists a cycle in the first container $C_{1}$, then the degree of $v_{0}$ will not double.
@@ 6 Back Arcs
Lemma 6.2 (No Back Arc to Early Container) effectively outlines a crucial restriction on back arcs in a container structured graph. It has effectively reduced this case of back arcs to one of interior degree doubling or back arcs mapping to the previous container.
@@
This was proved in the Lemmas 6.2 (No Back Arc to Early Container and 6.3 (First Back Arc Between Consecutive Containers) . ∎
@@ 9 Transitive Triangles and Seymour Diamonds
  • Case 1: If $d^{+}(y)=d$ and $d^{+}g(z)=d$: Both $y$ and $z$ have exactly $d$ outgoing connections. For the triangle to be transitive, they must not share all neighbors, meaning there exists at least one neighbor of $y$ that is not a neighbor of $z$. Thus, $N^{+}(y)\nsubseteq N^{++}(z)$.

### v2

@@ Abstract
Seymour’s Second Neighborhood Conjecture asserts that in the square of any oriented graph, there exists a node whose out-degree at least doubles. This paper presents a definitive proof of the conjecture by introducing the GLOVER (Graph Level Order) data structure, which facilitates a systematic partitioning of neighborhoods and an analysis of degree-doubling conditions. By leveraging this structure, we construct a decreasing sequence of subsets that establish a well-ordering of nodes, ensuring that no counterexample can exist. This approach not only confirms the conjecture for all oriented graphs but also provides a novel framework for analyzing degrees and arcs in complex networks. The findings have implications for theoretical graph studies and practical applications in network optimization and algorithm design.
@@ 1 Introduction
The problem of identifying structural properties in oriented graphs has been studied extensively, with particular focus on $m$-free graphs and Seymour vertices. Daamouch’s [7] work has been instrumental in this area, introducing significant results regarding induced subgraphs and their role in determining key vertex properties. This paper builds on these foundational ideas while presenting a novel approach that extends these concepts to interconnected container structures.
@@ 5 Theoretical Foundation
This definition of a container is conceptually related to Daamouch’s use of induced subgraphs in his study of $m$-free graphs [7]. However, the key difference lies in the formation of connected containers (see image), or neighboring containers, which allows for a more dynamic and scalable analysis of graph properties. This approach facilitates the identification of structural patterns across broader graph classes, a task that would otherwise be computationally prohibitive.
@@ A Graph Level Order (GLOVER) on set $S$ can be defined as follows:
Cycle in Container 1 If $G$ is an oriented graph represented as a GLOVER with $v_{0}$ as the chosen minimum degree node of degree $\delta$, and if there exists a cycle in the first container $C_{1}$ of length $\delta$, then the degree of $v_{0}$ will not double.
@@
The results presented in 5.1(Cycle in Container 1) and 5.5 (Exterior Less than Degree) are closely related to Daamouch’s [7] Lemma 1.1, which demonstrated that a minimum-degree vertex with no cycles in its induced graph has a degree that effectively doubles. While our work is similar to this result, we will extend the framework by leveraging container-based structures to provide a polynomial-time solution, highlighting the interplay between neighboring containers and their shared arcs.
@@
This next lemma looks into the Weak Decreasing Exteriors Lemma and considers the assumption of back arcs. There is a particular scenario that needs to be evaluated where a second neighbor of $y$ and first neighbor of $z$ can also be a first neighbor of x. The above lemma, called 5.6(Weak Decreasing Exteriors Lemma), has an explicit assumption that gets rid of this. In the next chapter we will do a lot of work to make sure that these back arcs are no longer a problem. However we will consider a second version of this lemma that gets rid of the back arcs assumption.
@@ 6 Back Arcs
Lemma 6.2 (No Back Arc to Early Container) effectively outlines a crucial restriction on back arcs in a container structured graph. It has effectively reduced this case of back arcs to one of interior degree doubling or back arcs mapping to the previous container. This is important because the previous section introduced the GLOVER framework as a total order of containers. Back arcs can disrupt this order. What we see next is a special case of how back arcs could disrput this order.
Lemma 6.3.
Back Arcs Special Case Let $G$ be an oriented graph represented by a GLOVER, with containers $C_{1},C_{2},\ldots,C_{n}$ defined based on the distance from the minimum degree node $v_{0}$ in order to prevent a degree doubling node.
1. Suppose $x\in C_{i}$ and $y\in C_{i+1}$, both satisfying the Decreasing Sequence Property (DSP). 2. The interior degree of $y$ with respect to $x$ doubles in $G[N^{+}(x)]^{2}$. 3. There exists a $w\in\text{ext}(x,y)$ such that the $int(y,w)$ doubles.
Then, for all $z\in\text{ext}(x,y)$, we have:
$|\text{ext}(y,z)|<|\text{ext}(x,y)|$
First we must notice that the order within the container system, which defines exterior arcs as going to the next container (i.e from container $C_{i}$ to container $C_{i+1}$) and not two containers down (i.e not from container $C_{i}$ to container $C_{i+2}$), explicitly forbids $x$ having an exterior neighbor $z$ in the container $C_{i+2}$ where $x$ is in container $C_{i}$. This gives the only possibility for $z$ as being an interior neighbor of $x$.
Consequently, the arc $w,z$ is an early back arc as we saw in 6.2(No Back Arc to Early Container). These arcs can lead to double counting. An out-neighbor of $w$ might be an out-neighbor of $x$, and a second out-neighbor of $y$. This implies that under the assumption that $w$’s interior degree double’s, $w$’s overall degree will double through that early back arc.
The framework under 6.2 (No Back Arc to Early Container) generalizes the solution to handle all potential back arc scenarios in the graph. Thus, we have strengthened the overall structure, making the algorithm robust against these issues and ensuring that the exterior degree conditions are preserved across all containers.
@@
This was proved in the Lemmas 6.2 (No Back Arc to Early Container and 6.4 (First Back Arc Between Consecutive Containers) . ∎
@@ 9 Transitive Triangles and Seymour Diamonds
  • Case 1: If $d^{+}(y)=d$ and $d^{+}(z)=d$: Both $y$ and $z$ have exactly $d$ outgoing connections. For the triangle to be transitive, they must not share all neighbors, meaning there exists at least one neighbor of $y$ that is not a neighbor of $z$. Thus, $N^{+}(y)\nsubseteq N^{++}(z)$.
@@ 10.2 Theorem Statement
This emergence of orientations of complete graphs as containers decrease in size adds a layer of symmetry and beauty to the Seymour Second Neighborhood Conjecture. There was never a question of these arising. Some questioned multiple nodes. Some questioned zero nodes. A complete graph occurrence though is a thing of beauty and it helps give rise to another result.
@@
The discovery of multiple degree-doubling nodes transforms the Seymour Second Neighborhood Conjecture from a question of existence to one of abundance. This result not only confirms the conjecture but reveals an inherent richness in the structure of oriented graphs. The presence of multiple nodes is tied to the emergence of complete graph structures as containers decrease in size, showcasing an element of mathematical beauty. Moreover, this has real-world applications: in network analysis, it suggests that influence or importance may not rest with a single node but instead within a clique or tightly connected group, offering new perspectives for studying collaborative dynamics and group influence.